# Zero-shot Medical Event Prediction Using a Generative Pre-trained Transformer on Electronic Health Records

### v1

@@
Objectives: Longitudinal data in electronic health records (EHRs) represent an individual‘s clinical history through a sequence of codified concepts, including diagnoses, procedures, medications, and laboratory tests. Foundational models, such as generative pre-trained transformers (GPT), can leverage this data to predict future events. While fine-tuning of these models enhances task-specific performance, it is costly, complex, and unsustainable for every target. We show that a foundation model trained on EHRs can perform predictive tasks in a zero-shot manner, eliminating the need for fine-tuning.
@@ Introduction
Recent work, CEHR-BERT (Pang et al., 2024), utilized a GPT-type model to generate synthetic EHR data. The model was pretrained on 3.7 million individuals’ medical histories, including conditions, medications, procedures, visit types, discharge facilities, and temporal data like start year, age, visit intervals, and durations. The GPT model was trained to learn the distribution of EHR sequences to generate new synthetic sequences. Predictive performance on synthetic data was evaluated using logistic regression on five tasks, such as HF readmission and hospitalization, showing synthetic data performed as well or better than real data. Similarly to the previous works, the model‘s zero-shot capabilities were never evaluated.
Task-specific fine-tuning of pre-trained models for a wide range of conditions and diseases becomes impractical due to high computational costs, challenges in cohort labeling, and the continuous need to update models with new clinical data. Therefore, a general model that could approach the performance of many fine-tuned models would simplify development, lower costs, and increase implementation, ultimately resulting in a greater impact on patient care. Zero-shot evaluation is defined as the ability of a model to make accurate predictions on unseen data without requiring additional training, thus allowing it to screen for multiple conditions and forecast a range of outcomes simultaneously. We present a zero-shot evaluation of a GPT model trained to forecast EHR concepts and examine the prospective performance of the model on various predictive tasks, including 12 specific diagnostic conditions and 14 broader categories of medical conditions.
@@ Methods
Our cohort comprises two decades of history from 2.4 million individuals who received care at UCLA Health sites across the greater Los Angeles region. To simulate a prospective deployment where the model would perform inference on data generated post-training, the model was trained using data before 2022 and tested with data acquired during and after 2022. The pre-2022 data set was randomly split into a training set (90%) and a validation set (10%). Data is structured according to the Observational Medical Outcomes Partnership (OMOP) Common Data Model (CDM). The OMOP CDM standardizes the format and content of EHR data, facilitating data integration, sharing, and analysis across diverse healthcare systems.
@@
As the EHR is temporally ordered, we utilize causal language modeling and build our framework on top of the GPT-2 architecture with the goal of forecasting future EHR entities. The main difference between traditional language modeling and the proposed approach is that the tokens $w_{i}$ represent medical events instead of words or subwords. For each individual $u_{i}$ defined as a sequence of tokens $u_{i}={w_{1},...w_{N}}$, where $|N|$ a vocabulary size, the objective of EHR modeling is similar to the language modeling:
$L(u_{i})=\sum_{j=1}^{N}logP(w_{i}|w_{i-1},w_{i-2},...,w_{0})$ (1)
Each individual is represented by a sequence of visits, separated by a separation token $s$, with each visit encompassing a time series of medical events, demographics, and special tokens that characterize the visit (see FIG. 1 a). Medical events include diagnostic $d$, medication $m$, procedure $p$, and laboratory codes $l$. The individual’s overall timeline begins with demographic information, such as gender and age. Each visit starts with a special time token that indicates the time difference relative to the previous visit. This time token, denoted as $t$, can take one of the following values: $t_{0}$ indicating the visit occurred within the same quarter as the previous visit; $t_{1}$ signifying a visit that occurred more than three months but within six months of the previous visit; $t_{2}$ denotes a visit that occurred later than six months but within a year; and $t_{3}$ represents a visit that occurred more than a year later. The time token is followed by visit-type token $vt$ at the beginning of each visit. Each visit can end with discharge type tokens $dt$.
We trained our model using the common next-token prediction objective used in pretraining language models.
Zero-shot evaluation cohort
@@
The model‘s forecasting performance is evaluated using precision and recall metrics on the visit level. Precision is defined as $\frac{TP}{TP+FP}$ and recall is defined as $\frac{TP}{TP+FN}$, where TP is the number of true positive tokens, FN - false negative tokens, FP - false positive tokens. During inference, given a sequence of previous visits, the model predicts the next visit by forecasting subsequent concepts along the timeline until a separation token is predicted. The top 1 greedy decoding path was considered at each time step. Precision and recall are computed between the predicted sequence and the ground truth visit.
@@ Results
The evaluation results of the model’s zero-shot capabilities for 12 diagnostic tasks at each prediction window (three months and six months) are presented in FIG.2,S1 and for 14 higher-level diagnostic tasks in FIG.3,S2. The results demonstrate the model‘s capability to predict multiple disease outcomes with varying levels of accuracy depending on the disease, prediction window, and the number of forecasts $N$ (e.g., top 1, top 5, top 10, or top 20). Our analysis reveals that the model’s predictive performance varies across different diseases and time horizons. As we expand the number of top predictions considered (from top 1 to top 20), we observe a consistent increase in TP across all conditions as it allows the model to evaluate a wider range of possible outcomes at each time step. However, a similar trend is observed for false positives, which also tend to rise as $N$ increases. For example, the model predicted that 9.37% of individuals diagnosed with chronic kidney disease would develop the condition within the three-month prediction window for $N$=1. When $N$ was increased to 20, the percentage of TP rose to 84.1%, accompanied by a corresponding rise in FP to 68.47% within the same three-month window. The same trend was observed for individuals diagnosed with chronic obstructive pulmonary disease. While 3.54% of individuals diagnosed with the condition were predicted to be positive by the model for $N$=1, the percentage increased to 54.89% as $N$ increased to 20. The model could predict 4.43% TP individuals with diseases from the ’C00-D49 neoplasms’ category at $N$=1, with this value increasing to 79.12% as $N$ increased to 20. A corresponding increase was observed in the number of FP, rising from 1.68% at $N$=1 to 56.78% at $N$=20. Generally, the 3-month predictions show higher TP and less FP than the 6-month predictions across all conditions. An example that highlights this trend is the model’s performance on heart failure, predicting that 5.73% of individuals would experience heart failure within the three-month prediction window and 4.57% within the six-month window for $N$ equal to 1. Only 0.16% of individuals were FP within the three-month prediction window, while this figure increased slightly to 0.23% for the six-month window. When $N=20$, forecasts at each time step lead to an increase in the number of TP individuals up to 53.6% with a corresponding increase in FP up to 10.29% within the three-month prediction window. The number of TP dropped slightly to 47.97% for the six-month window, with a slight increase in FP to 12.98%.
@@ Discussion
Many EHR codes are highly correlated due to the interconnected nature of medical conditions and treatments. Selecting a single code for fine-tuning or zero-shot evaluation can bias results and overestimate model performance. The model may rely on correlations with related codes in an individual’s history rather than demonstrating true causal understanding. For instance, major depressive disorder codes (F32, F33, F34) often follow anxiety disorders (F41), mood disorders (F39), or personality disorders (F60). Similarly, heart failure (I50) is frequently preceded by hypertensive heart disease (I11, I13) or cardiomyopathy (I42).
@@ Conclusion
Supplementary Material

### v2

@@
Objectives: Longitudinal data in electronic health records (EHRs) represent an individual‘s clinical history through a sequence of codified concepts, including diagnoses, procedures, medications, and laboratory tests. Generative pre-trained transformers (GPT) can leverage this data to predict future events. While fine-tuning of these models can enhance task-specific performance, it becomes costly when applied to many clinical prediction tasks. In contrast, a pretrained foundation model can be used in zero-shot forecasting setting, offering a scalable alternative to fine-tuning separate models for each outcome.
@@ Introduction
Recent work, CEHR-BERT (Pang et al., 2024), utilized a GPT-type model to generate synthetic EHR data. The model was pretrained on 3.7 million individuals’ medical histories, including conditions, medications, procedures, visit types, discharge facilities, and temporal data like start year, age, visit intervals, and durations. The GPT model was trained to learn the distribution of EHR sequences to generate new synthetic sequences. Predictive performance on synthetic data was evaluated using logistic regression on five tasks, such as HF readmission and hospitalization, showing that synthetic data performed as well or better than real data. Similarly to the previous works, the model‘s zero-shot capabilities were never evaluated.
Task-specific fine-tuning of pre-trained models for a wide range of conditions and diseases becomes impractical due to high computational costs, challenges in cohort labeling, and the continuous need to update models with new clinical data. Therefore, a general model that could approach the performance of many fine-tuned models would simplify development, lower costs, and increase implementation, ultimately resulting in a greater impact on patient care. Zero-shot evaluation is defined as the ability of a model to make accurate predictions on unseen data without requiring additional training, thus allowing it to screen for multiple conditions and forecast a range of outcomes simultaneously. We present a zero-shot evaluation of a GPT model trained to forecast EHR concepts and examine the prospective performance of the model on various predictive tasks, including 12 specific diagnostic conditions and 14 broader categories of medical conditions. While our focus is on zero-shot inference, we also report results from task-specific fine-tuning to benchmark potential gains from full supervision.
@@ Methods
Our cohort comprises two decades of history from 2.4 million individuals who received care at UCLA Health sites across the greater Los Angeles region (see TABLE 1). To simulate a prospective deployment where the model would perform inference on data generated post-training, the model was trained using data before 2022 and tested with data acquired during and after 2022. The pre-2022 data set was randomly split into a training set (90%) and a validation set (10%) at the patient level. Data is structured according to the Observational Medical Outcomes Partnership (OMOP) Common Data Model (CDM). The OMOP CDM standardizes the format and content of EHR data, facilitating data integration, sharing, and analysis across diverse healthcare systems.
Table 1: Demographic characteristics of the study cohort.
Characteristic Value N (%)
Age (years) Mean (SD) 53.3 (20.0)
Sex
 Male 820 49.2%
 Female 847 50.8%
Race
 White 1001400 44.9%
 Black or African American 99894 4.7%
 Asian 180447 8.1%
 Native Hawaiian or Other Pacific Islander 4665 0.3%
 Multirace 85774 3.8%
 Other / Unknown 849267 38.2%
@@
For each individual $u_{i}$ the full EHR sequence is organized as:
$u_{i}=(V_{1},\texttt{[SEP]},V_{2},\texttt{[SEP]},\ldots,V_{L}),$ (1)
where each visit $V_{t}=(w_{t,1},w_{t,2},\ldots,w_{t,m_{t}})$ consists of $m_{t}$ medical events, demographics, and special tokens that characterize the visit (see FIG. 1 a). Medical events include diagnostic $d$, medication $m$, procedure $p$, and laboratory codes $l$. The individual’s overall timeline begins with demographic information, such as gender and age. Each visit starts with a special time token that indicates the time difference relative to the previous visit. This time token, denoted as $t$, can take one of the following values to indicate when the visit occurred relative to the previous visit: $t_{0}$, visit occurred within the same quarter; $t_{1}$, visit occurred more than three months but within six months; $t_{2}$, visit occurred later than six months but within a year; and $t_{3}$, visit occurred more than a year later. The time token is followed by a visit-type token $vt$ at the beginning of each visit. Each visit can end with discharge type tokens $dt$. Each event token $w_{v,j}$ is associated with a timestamp $\tau_{v,j}$ such that $\tau_{v,j}\leq\tau_{v,j+1}$ within the same visit. The objective of EHR modeling is then similar to language modeling:
$\mathcal{L}(u_{i})=\sum_{t=1}^{L}\sum_{j=1}^{m_{t}}\log P(w_{t,j}\mid w_{1,1},\ldots,w_{t,j-1})$ (2)
Evaluation cohort
@@
The model’s forecasting performance is evaluated at the visit level. For a given patient sequence, the first visit is used as input to generate the second visit, and the predicted visit is then compared to the actual second visit. Next, the first two observed visits are provided as input to generate the third visit, and so on. This stepwise evaluation continues through the entire timeline, ensuring that each prediction is made using only prior information. At each step, the model autoregressively generates a sequence of medical concepts—such as diagnoses, procedures, medications, and labs—until a separation token is produced, indicating the end of the visit. We use greedy decoding (top-1 selection) at each time step to produce the output tokens. Precision and recall are computed by comparing the set of predicted tokens for the visit against the corresponding ground truth tokens, where precision is defined as $\frac{TP}{TP+FP}$ and recall as $\frac{TP}{TP+FN}$.
@@
Fine-tuning Setup for Diagnostic Prediction
To compare zero-shot forecasting with supervised learning, we conducted a fine-tuning experiment using a binary classification task for each evaluation cohort. We used the pretrained GPT model as a frozen feature extractor. The model’s output embeddings were pooled using mean pooling across tokens and passed through a dropout layer, followed by a linear classification head trained for binary prediction. Only the classification head was updated during training, while the transformer weights remained frozen. The model was trained using the Adam optimizer with a learning rate of 1e-4 with class balancing via weighted sampling. We measured the number of true positives (TP) achieved at the same number of false positives (FP) as observed in the results of zero-shot evaluation for the same disease code. This allowed direct comparison of classification utility between fine-tuning and zero-shot evaluation under equal specificity constraints.
@@ Results
The evaluation results of the model’s zero-shot capabilities for 12 diagnostic tasks at each prediction window (three months and six months) are presented in FIG.2,S2 and for 14 higher-level diagnostic tasks in FIG.3,S3. The results demonstrate the model‘s capability to predict multiple disease outcomes with varying levels of accuracy depending on the disease, prediction window, and the number of forecasts $N$ (e.g., top 1, top 5, top 10, or top 20). Our analysis reveals that the model’s predictive performance varies across different diseases and time horizons. As we expand the number of top predictions considered (from top 1 to top 20), we observe a consistent increase in TP across all conditions as it allows the model to evaluate a wider range of possible outcomes at each time step. However, a similar trend is observed for false positives, which also tend to rise as $N$ increases. For example, the model predicted that 9.37% of individuals diagnosed with chronic kidney disease would develop the condition within the three-month prediction window for $N$=1. When $N$ was increased to 20, the percentage of TP rose to 84.1%, accompanied by a corresponding rise in FP to 68.47% within the same three-month window. The same trend was observed for individuals diagnosed with chronic obstructive pulmonary disease. While 3.54% of individuals diagnosed with the condition were predicted to be positive by the model for $N$=1, the percentage increased to 54.89% as $N$ increased to 20. The model could predict 4.43% TP individuals with diseases from the ’C00-D49 neoplasms’ category at $N$=1, with this value increasing to 79.12% as $N$ increased to 20. A corresponding increase was observed in the number of FP, rising from 1.68% at $N$=1 to 56.78% at $N$=20. Generally, the 3-month predictions show higher TP and less FP than the 6-month predictions across all conditions. An example that highlights this trend is the model’s performance on heart failure, predicting that 5.73% of individuals would experience heart failure within the three-month prediction window and 4.57% within the six-month window for $N$ equal to 1. Only 0.16% of individuals were FP within the three-month prediction window, while this figure increased slightly to 0.23% for the six-month window. When $N=20$, forecasts at each time step lead to an increase in the number of TP individuals up to 53.6% with a corresponding increase in FP up to 10.29% within the three-month prediction window. The number of TP dropped slightly to 47.97% for the six-month window, with a slight increase in FP to 12.98%.
@@
Fine-tuning evaluation results
Table 2 reports the number of true positives (TP) achieved via fine-tuning (TPf) and in the zero-shot setting (TPz) across 12 diagnostic tasks, evaluated at a fixed number of false positives (FP). Overall, while fine-tuning provides performance gains, the absolute differences in TP between TPf and TPz are generally modest. For instance, in the 3-month prediction window, the average TP improvement from fine-tuning across tasks is 3.3, with particularly noticeable gains for diagnoses like I25 (chronic ischemic heart disease: 19.62 vs. 11.84) and J44 (chronic obstructive pulmonary disease: 7.01 vs. 2.67). However, in several other tasks—such as M32 (systemic lupus erythematosus) and C22 (liver cancer)—zero-shot performance is within 0.5–1.0 TP of the fine-tuned counterpart. Similar trends are observed at the 6-month horizon. Fine-tuning yields higher TPs in most tasks, but the zero-shot model often remains competitive. Notably, in tasks with low prevalence or more complex diagnostic criteria (e.g., C25, pancreatic cancer), both settings achieve low TPs, indicating shared difficulty rather than fine-tuning efficacy. Overall, these results suggest the robustness of the zero-shot representation. Furthermore, these performance deltas must be interpreted in light of the resource demands associated with extensive hyperparameter tuning for task-specific fine-tuning. In contrast, the zero-shot model operates without task-specific training, making it highly scalable across new tasks.
I50 N18 E11 C22 C71 C25 J44 M06 M32 I25 I63 F32
3 months
TPf 6.86 17.2 13.86 2.94 2.51 0.00 7.01 1.87 2.44 19.62 2.97 9.64
TPz 4.57 9.37 10.50 3.53 2.35 0.00 2.67 2.17 2.93 11.84 0.77 5.41
FP 0.23 1.19 1.39 0.00 0.00 0.00 0.28 0.07 0.06 2.03 0.12 2.01
6 months
TPf 6.86 15.05 14.4 1.03 2.47 0.00 4.87 1.29 1.45 15.12 2.42 11.18
TPz 4.57 8.40 8.32 1.47 2.13 0.00 2.67 2.15 2.04 10.03 0.77 5.41
$FP$ 0.23 1.44 1.63 0.00 0.00 0.00 0.36 0.11 0.09 2.34 0.18 2.10
Table 2: TPf denotes the true positives achieved through fine-tuning, while TPz represents the true positives obtained in the zero-shot setting, both evaluated at a fixed number of false positives across 12 different diagnostic tasks.
@@ Discussion
Fine-tuning increased the percentage of correctly identified cases for certain diagnostic conditions, in other cases, the zero-shot model already performed competitively. This suggests that while fine-tuning can provide marginal gains, its benefits are not uniform across tasks and may depend on the nature and prevalence of the target condition. Importantly, our fine-tuning setup did not involve full hyperparameter optimization; instead, we used a fixed training configuration to ensure a consistent and practical approach. While performance could likely be further improved with tuning, in our experience, such improvements would be modest and not change the overall interpretation of our results. Moreover, extensive tuning would increase the computational cost.
@@
Many EHR codes are highly correlated due to the interconnected nature of medical conditions and treatments. Selecting a single code for fine-tuning or zero-shot evaluation can bias results and overestimate model performance. The model may rely on correlations with related codes in an individual’s history rather than demonstrating true causal understanding. For instance, major depressive disorder codes (F32, F33, F34) often follow anxiety disorders (F41), mood disorders (F39), or personality disorders (F60). Similarly, heart failure (I50) is frequently preceded by hypertensive heart disease (I11, I13) or cardiomyopathy (I42). To better understand these patterns, we performed an error analysis, with detailed results provided in the Supplementary Materials (Figure S1).
@@ Conclusion
Supplementary Materials
Model Architecture and training configuration
We pretrained a causal language model on longitudinal EHR sequences using a modified GPT-2 architecture implemented with the HuggingFace Transformers library. The configuration and training setup are summarized below.
Model Architecture.
The architecture details are provided in Table S3.
Parameter Value
Architecture GPT2LMHeadModel
Hidden size 384
Number of layers 4
Attention heads 8
Context window 513 tokens
Vocabulary size 4840
BOS/EOS token ID 4839
Total parameters $\sim$56 million
Table S3: GPT model configuration used for EHR sequence modeling.
Training Setup.
The model was trained on tokenized patient timelines truncated to 512 tokens per sequence. Each input takes the form:
$u_{i}=(\texttt{[CLS]},V_{1},\texttt{[SEP]},V_{2},\texttt{[SEP]},\ldots,V_{L}),$
where each visit $V_{t}=(w_{t,1},w_{t,2},\ldots,w_{t,m_{t}})$ includes diagnosis, medication, procedure, and laboratory tokens.
  • Batch size: 256
  • Optimizer: AdamW
  • Learning rate: 3e-4
  • Epsilon: 1e-8
  • Training epochs: 2
  • Precision: Full FP32
Loss Function.
The model was trained using standard causal language modeling loss (next-token prediction), with attention masking applied to ignore padded positions.
Reproducibility.
Model checkpoints and training losses were saved per epoch. The data was split at the patient level to ensure no leakage between training, validation, and test sets.
Error analysis
To further explore the behavior of the pretrained GPT model, we visualized patient-level embeddings using UMAP for several diagnostic conditions. For each condition, we extracted the final-layer embeddings generated by the model for individual patients and projected them into a two-dimensional space using UMAP. We then labeled each point as a TP or FP based on the model’s zero-shot prediction outcome. As shown in Figure S1, there is substantial overlap between TP and FP patient representations across multiple diseases. This overlap suggests that the model assigns high confidence to patients who are, in fact, clinically similar to true positives, but ultimately do not receive the diagnosis. These errors may stem from limitations in the model’s ability to distinguish subtle clinical differences or from ambiguity in the diagnostic labels themselves (e.g., due to underdiagnosis, delayed documentation, or overlapping comorbidities). Future work is needed to better understand these borderline cases and whether incorporating richer supervision signals or patient-level context could reduce such errors.
Figure S1: UMAP visualization of patient embeddings generated by the pretrained GPT model for nine diagnostic conditions. Each point represents an individual patient, colored by prediction outcome (true positive or false positive).